# Robust Adaptive Time-Varying Control Barrier Function with Application to Robotic Surface Treatment

### v1

@@ V-B Scenario Setup for Simulation and Experiment
In this work, we focus on the quality guarantees in the dotted box as shown in Fig. 3, rather than how to obtain MRR from the user and estimate the contact area. It is worth noting that obtaining MRR and contact area estimation can be done with analytical and numerical approaches [23][24]. We arbitrarily select the desired MRR and design a simple but reliable plane-based scenario as shown in the middle of Fig. 3, so that the contact area can be easily calculated with the closed solution between the circle polishing tool and the plane. This allows us to show a clear comparison across CBFs-based controllers in the simulation and real robotic experiment. We first let a robot polish on the rectangular aluminium plane and pass edge areas and return to the start point after moving around the plane as shown in the middle of Fig. 3. On the edge regions, the contact areas are changed, which means that the current contact forces should be also adjusted to let the current MRR remain in the lower and upper MRR bounds (e.g. at ±$\pm 10$ of the desired MRR here). To this end, we calculate $\bm{f}_{\mathsf{lower}}$ and $\bm{f}_{\mathsf{upper}}$ from the lower/upper MRR bounds by using (1), which is shown in Fig. 4(a). In this scenario, we use a P controller as the nominal controller in the simulation. Note that we intentionally define the desired contact force profile for the nominal controller outside of the force bound, (28) to clearly show whether each controller ensures the bounds in the presence of uncertainties.
@@ V-C Simulation Results
Fig. 4 shows the simulation results of controllers in the robotic surface treatment scenario. We first carry out Time-Varying CBF (TVCBF) [15] as a baseline in the simulation. In this simulation, we set the ground truth of $\bm{k},\bm{b}$ to $1400$ and $70$, respectively, on the other hand, the system model has the putative parameters, $\hat{\bm{k}},\hat{\bm{b}}$ that are $1000$ and $10$, respectively. Note that the proposed method does not rely on the initial estimate conditions such as over or under estimate of the parameters, as long as the initial conditions satisfy the maximum possible error condition. To show a clear comparison between controllers, we use the same $\mathcal{K}_{\infty}$ functions as $K(\cdot)$ and $\alpha(\cdot)$ in (4) and (25), respectively. As shown in Fig. 4(a) and Fig. 4(b), TVCBF violates the constraint of force bound, (28) due to the uncertainties of model and input disturbances. On the other hand, the proposed method, RaTVCBF complies with the constraint, $h\geq 0$ since we impose the robustness against the parametric model uncertainty and input disturbances by using the maximum possible error, $\tilde{\bm{\vartheta}}$ and the bound of the disturbances, $\delta$, respectively. Since the method ensures that the contact forces stay within the force bounds, we can expect that it also enables the current MRR to remain within the desired lower and upper bounds of the desired MRR while the baseline violates them as shown in Fig. 4(c). Fig. 4(b) and 4(c) show the performances of the proposed method with and without SMID. It is observed that applying SMID to RaTVCBF reduces the conservatism by approximately $48\%$ based on each mean value. We use the batch size of $5$ and $\varepsilon=0.0008$ in this simulation.
@@ V-D Experimental Validation
We first generate the robot trajectory with constant linear velocity of the tool and calculate the contact area based on the area of circular segment in edge regions. Fig. 5 shows the performances of each controller. The proposed method can regulate contact forces to stay within the force bounds compared to TVCBF as shown in Fig. 5(a). The values of $h$ functions as shown in Fig. 5(b) indicate clearly whether each method violates the force bounds. Since the proposed method allows the contact forces to maintain within (28), the proposed method ensures the guaranteed range of MRR while the baseline violates the range as shown in Fig. 5(c). This is because we enforce robustness against uncertainties by using the maximum possible error and the bound of input disturbances. Moreover, we use SMID to reduce the conservative behaviors of RaTVCBF based on the collected input-output data as shown in Fig. 5(b) and 5(c). It is observed that leveraging SMID alleviates the conservatism by about $30\%$ when comparing each mean value. We have the batch size of $5$ and $\varepsilon=0.35$ in this experiment. For practical implementation, $\varepsilon$ is a tunable parameter determined by the magnitude of disturbances, measurement noise, and unmodeled dynamics [20]. In addition, $k_{p}$ from (1) is an unknown parameter, but it can be encapsulated in the parametric structure of $\theta$, (e.g. $\theta$ =$\begin{bmatrix}k_{p}k&k_{p}b\end{bmatrix}^{\top}$), so this uncertainty can also be treated in the proposed method. In the presence of tool wear, the Preston coefficient $k_{p}$ will be time-varying and may be estimated based on the ratio between initial and final MRR [25].

### v2

@@ V-B Scenario Setup for Simulation and Experiment
In this work, we focus on the quality guarantees in the dotted box as shown in Fig. 3, rather than how to obtain MRR from the user and estimate the contact area. It is worth noting that obtaining MRR and contact area estimation can be done with analytical and numerical approaches [XIAO2020188][YkIROS2022]. We arbitrarily select the desired MRR and design a simple but reliable plane-based scenario as shown in the middle of Fig. 3, so that the contact area can be easily calculated with the closed solution between the circle polishing tool and the plane. This allows us to show a clear comparison across CBFs-based controllers in the simulation and real robotic experiment. We first let a robot polish on the rectangular aluminium plane and pass edge areas and return to the start point after moving around the plane as shown in the middle of Fig. 3. On the edge regions, the contact areas are changed, which means that the current contact forces should be also adjusted to let the current MRR remain in the lower and upper MRR bounds (e.g. at $\pm 10$ of the desired MRR here). To this end, we calculate $\bm{f}_{\mathsf{lower}}$ and $\bm{f}_{\mathsf{upper}}$ from the lower/upper MRR bounds by using (1), which is shown in Fig. 4(a). In this scenario, we use a P controller as the nominal controller in the simulation. Note that we intentionally define the desired contact force profile for the nominal controller outside of the force bound, (28) to clearly show whether each controller ensures the bounds in the presence of uncertainties.
@@ V-C Simulation Results
Fig. 4 shows the simulation results of controllers in the robotic surface treatment scenario. We first carry out Time-Varying CBF (TVCBF) [TvCBF2019] as a baseline in the simulation. In this simulation, we set the ground truth of $\bm{k},\bm{b}$ to $1400$ and $70$, respectively, on the other hand, the system model has the putative parameters, $\hat{\bm{k}},\hat{\bm{b}}$ that are $1000$ and $10$, respectively. Note that the proposed method does not rely on the initial estimate conditions such as over or under estimate of the parameters, as long as the initial conditions satisfy the maximum possible error condition. To show a clear comparison between controllers, we use the same $\mathcal{K}_{\infty}$ functions as $K(\cdot)$ and $\alpha(\cdot)$ in (4) and (25), respectively. As shown in Fig. 4(a) and Fig. 4(b), TVCBF violates the constraint of force bound, (28) due to the uncertainties of model and input disturbances. On the other hand, the proposed method, RaTVCBF complies with the constraint, $h\geq 0$ since we impose the robustness against the parametric model uncertainty and input disturbances by using the maximum possible error, $\tilde{\bm{\vartheta}}$ and the bound of the disturbances, $\delta$, respectively. Since the method ensures that the contact forces stay within the force bounds, we can expect that it also enables the current MRR to remain within the desired lower and upper bounds of the desired MRR while the baseline violates them as shown in Fig. 4(c). Fig. 4(b) and 4(c) show the performances of the proposed method with and without SMID. It is observed that applying SMID to RaTVCBF reduces the conservatism by approximately $48\%$ based on each mean value. We use the batch size of $5$ and $\varepsilon=0.0008$ in this simulation.
@@ V-D Experimental Validation
We first generate the robot trajectory with constant linear velocity of the tool and calculate the contact area based on the area of circular segment in edge regions. Fig. 5 shows the performances of each controller. The proposed method can regulate contact forces to stay within the force bounds compared to TVCBF as shown in Fig. 5(a). The values of $h$ functions as shown in Fig. 5(b) indicate clearly whether each method violates the force bounds. Since the proposed method allows the contact forces to maintain within (28), the proposed method ensures the guaranteed range of MRR while the baseline violates the range as shown in Fig. 5(c). This is because we enforce robustness against uncertainties by using the maximum possible error and the bound of input disturbances. Moreover, we use SMID to reduce the conservative behaviors of RaTVCBF based on the collected input-output data as shown in Fig. 5(b) and 5(c). It is observed that leveraging SMID alleviates the conservatism by about $30\%$ when comparing each mean value. We have the batch size of $5$ and $\varepsilon=0.35$ in this experiment. For practical implementation, $\varepsilon$ is a tunable parameter determined by the magnitude of disturbances, measurement noise, and unmodeled dynamics [smidbook]. In addition, $k_{p}$ from (1) is an unknown parameter, but it can be encapsulated in the parametric structure of $\theta$, (e.g. $\theta$ $\begin{bmatrix}k_{p}k&k_{p}b\end{bmatrix}^{\top}$), so this uncertainty can also be treated in the proposed method. In the presence of tool wear, the Preston coefficient $k_{p}$ will be time-varying and may be estimated based on the ratio between initial and final MRR [Quinsat].